a spherical portion has been removed from a solid sphere having a charge distributed uniformly in its volume. the electric field inside the emptied space is non-zero and uniform.how this can be  proved using superposition principle? 

Grade:Upto college level

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Nandan

assume in the cavity charge density p and -p.

let position of point from complete sphere is R vector and from cavity centre is r vector

net elec field= pR/3εo - pr/3εo = p(R-r)/3εo

= pa/3εo (from vector addition)   

where a vector is from centre of sphere to centre of cavity. so we take any point inside the cavity elec field will be pa/3ε and it is independent of position of point in the cavity. hence uniform.

 

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively :Click here to download the toolbar..